### v1

# skipabove=0.5skipbelow=0.5align=center

[…]
skipabove=0.5skipbelow=0.5align=center
Conceptualizing Security Reputation in the Software Supply Chain: A Case for Actor-Based Metrics
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Conceptualizing Security Reputation in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
A Case for Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Track Bad Engineers, Not Just Bad Code: Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
A Case for Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
AMSEC: Actor-based metrics for Software Supply Chain Security
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Repitor: A Case for Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
ARMS: Actor Reputation Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
ARMS: A Vision for Actor Reputation Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
ARMS: A Vision for Actor Reputation Metric Systems in the Open-Source Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Abstract
Many critical information technology and cyber-physical systems rely on a supply chain of open-source software projects. OSS project maintainers often integrate contributions from external actors. While maintainers can assess the correctness of a change request, assessing a change request’s cybersecurity implications is challenging. To help maintainers make this decision, we propose that the open-source ecosystem should incorporate Actor Reputation Metrics (ARMS). This capability would enable OSS maintainers to assess a prospective contributor’s cybersecurity reputation. To support the future instantiation of ARMS, we identify seven generic security signals from industry standards; map concrete metrics from prior work and available security tools, describe study designs to refine and assess the utility of ARMS, and finally weigh its pros and cons.
Index Terms:
Software Supply Chain, Reputation systems
I Introduction
Most commercial software depends on open-source software components [1]. Although this approach reduces development costs, it results in a software supply chain that exposes an organization to cybersecurity risks [2, 3]. Many prior works have examined software supply chain security failures and have developed security techniques [4, 5, 6], engineering processes and frameworks [7, 8, 9]. These works have had substantial success, e.g., the now widely-used Sigstore project for provenance [4], and OpenSSF Scorecard project [10] for process. However, prior works have paid little attention to the actor element of the software supply chain [8].
In this vision paper, we propose ARMS, an Actor Reputation Metric System, to track the security qualifications of engineers in the open-source software supply chain. We first define ARMS’s requirements based on the threat model in this context. We then propose a conceptual design for a reputation-based framework that evaluates an actor’s trustworthiness. Next, to obtain indicators of security skill and expertise, we map high-level recommendations from frameworks like SLSA and CNCF to specific, measurable metrics derived from prior research and existing security tools. We outline evaluations to assess the implementation and effectiveness of an ARMS system. Finally, we discuss potential future directions and improvements for our approach.
Our proposal explores the development and operationalization of actor-based metrics to address software supply chain security failures. While our proposal requires careful implementation to align with developers’ perspectives and project needs, we hope it brings greater research attention to the actor side of the software supply chain.
Our contributions are:
  • We propose Actor Reputation Metric Systems (ARMS) to support open-source maintainers in vetting contributions from unknown engineers. Focusing on cybersecurity, we identify signals that could indicate cybersecurity expertise, and metrics that could be used to operationalize those signals.
  • We design experiments that could be used to evaluate ARMS, and discuss the pros and cons of deploying such a system.
II Background and Motivation
II-A The Open-Source Software Supply Chain
Open-source software is widely integrated into commercial [11] and government [12] systems. Any individual open-source component is developed by a maintainer team. With their approval, outsiders may be permitted to contribute code [13]. Beyond this direct incorporation of external contributions, each such project often depends on others as components, recursively. This web of interdependencies is a feature of open-source development, allowing (in an idealized world) a reduction in repeated effort [14]. However, each additional point of trust increases the potential attack surface. From the perspective of the downstream application, the result is a software supply chain that can be attacked either through its artifacts or through its actors [15].
We follow the software supply chain definition of Okafor et al. [8]: in their production and distribution, software artifacts undergo a series of operations overseen by actors. This definition indicates that a software supply chain can be secured only through attention to all of these entities.
II-B Artifact-Based Evaluations Are Not Enough
A key activity in open-source projects is expanding the actor pool by introducing new maintainers and contributors into projects [13]. As a software package gains popularity, interest from potential contributors increases [16, 17, 18], often resulting in onboarding new maintainers and merging change requests from new contributors. Evaluating these individuals may involve reputational factors such as community status and connections [19, 20, 16], but security considerations are rarely enforced due to the challenges OSS teams face in managing security resources effectively [21, 22].
As argued by Okafor et al. [8], most cybersecurity work takes an artifact‐based perspective. Efforts to develop static and dynamic security analysis tools (SAST, DAST) [23], refine code reviews [24, 25], and assess artifact provenance [26, 4, 27] are all focused on confirming that an artifact is, to the limits provided by the technique, reliable. While this is not an unreasonable strategy, there are many reasons to avoid relying exclusively on artifact checks, such as:
  • Reliance on known vulnerabilities. Automated scanners typically match code against vulnerability databases; therefore, zero‑day flaws or novel attack vectors often evade detection[28] e.g., Log4j [29].
  • Biased human review. Manual code reviews and audits are applied inconsistently, frequently favoring contributors familiar to project maintainers [19, 18, 30].
  • Scalability constraints. As projects grow, sustaining thorough artifact checks becomes resource‑intensive, leading to superficial reviews or delayed patching [30, 31].
  • Limited socio‑technical insight. Artifact‑centric methods inspect code but ignore the developer behaviors and workflows that often precipitate security issues [32].
These shortcomings underscore the need for actor-based security measures, rather than considering code in isolation from its author. We thus turn now to reputation as a means of estimating the quality of an actor’s contributions.
II-C Using Actor Reputation to Establish Trust
Reputation systems establish trust between parties who have not been previously connected [33]. When social systems integrate reputation, incentives associated with positive reputation can encourage good behavior over time [34]. Following Hendrikx et al. [35], any reputation system has three interacting entities:
  • Trustor: The party placing trust (e.g., the maintainer team).
  • Trustee: The party being evaluated (e.g., a potential contributor).
  • Trust Engine (Recommender): The broker that supplies the trustor with information about the trustee (interaction data). Its design varies by application and threat model — e.g., for communication [33], online-auction [36], etc.
Two common examples of reputation systems in software engineering are GitHub’s star system [37] and the Stack Overflow point-based system [38]. Both of these systems can be used by a trustor to quantify the number of users satisfied with a trustee’s projects and contributions [39]. As a result, they might influence which GitHub projects an engineer (trustor) may deem to be reliable [40], and which Stack Overflow answers may be trusted by their readers [41].
In the following sections, we introduce ARMS to formalize the concepts of an actor reputation metric system to promote cybersecurity within the OSS ecosystem.
III Threat Model
III-A Threat Actors
The goal of ARMS is to provide project maintainers with measurements of the security expertise of prospective contributors or maintainers. ARMS operates within a context with three kinds of threat actors:
  1. 1.
    Inexperienced Contributors/Inadvertent Vulnerability: Contributors who lack sufficient security expertise — e.g., they are unfamiliar with standard security practices and tooling within the ecosystem — attempt to join or maintain OSS projects, potentially introducing vulnerabilities through mistakes [42, 43].
  2. 2.
    Reputation Spoofing: Malicious actors deliberately craft the appearance of security expertise to gain collaborator or maintainer status.
  3. 3.
    Impersonation. Impersonation occurs when a malicious actor gains control of a legitimate user’s account (e.g., via key compromise [44, 45]).
Our ARMS approach considers the first two classes of threat. The third class, impersonation threats, are out of scope — they undermine the assumption of stable identities necessary for a reputation-based system [46].
III-B Examples
We give examples of each kind of threat we outlined above.
III-B1 Dexcom (Inadvertent Vulnerability)
Dexcom is a medical device company whose products include continuous glucose monitors (CGMs) used by diabetics [47]. Their CGM products were the first to incorporate “smart” capabilities such as pushing health notifications to one’s smartphone. In 2019, Dexcom’s engineers made an error leading to a service outage, resulting in a lack of notifications; many were hospitalized and at least one death is attributable [48]. This was the second such outage in a 12-month window. Although we presume that Dexcom is not intentionally harming its customers, its engineers’ inability to sustain a safety-critical system suggests inadequate experience for this class of work.
III-B2 XZ Utils Backdoor (Reputation Spoofing)
In March 2024, a backdoor was discovered in XZ Utils, an open-source compression tool, which allowed attackers to gain root privileges and execute malicious code on affected systems [49]. The vulnerability was introduced by an actor who had built trust within the project through non-malicious contributions, and they were eventually promoted to co-maintainer [49]. In retrospect, this particular actor had several suspicious reputational signals, most notably having no accounts on other sites and no history of contributions to other projects.
III-B3 ESLint Credential Compromise (Impersonation)
ESLint, a widely used static analysis tool in the npm ecosystem for scanning JavaScript code, was compromised on July 12, 2018 [45]. Attackers gained access to a maintainer’s npm credentials and published malicious package updates to the npm registry [50]. Because the attacker used the identify of a reputable maintainer, a reputation system (which assume stable identities) could not anticipate this attack.
Figure 1: Overview of proposed ARMS system and context case study. Potential contributors (trustees), who may be malicious (Actor A), inadequately expertised (Actor B), or genuine (and capable) (Actor C), express interest or submit change requests . The maintainer team (trustors) requests reputation information on these contributors. The ARMS system retrieves each contributor’s interaction history and quantifies it using the defined security signals and metrics (Interaction data formatter & security signal scoring). Next, the reputation calculator weights these signal values by package usage, community tenure, and centrality, then composites the results and compares them to ecosystem-wide benchmarks (Impact & Benchmark Scoring). Finally, each trustee’s reputation score and recommended action are provided to the maintainer team.
IV ARMS Conceptual Model
To formalize an OSS actor reputation system for GitHub, we propose a reputation system based on the reference model of Hendrikz et al. [35] described earlier, and adapt it to the OSS supply chain context. The following subsections outline our proposed system, define security metrics for an actor’s security reputation, and propose evaluation metrics to measure the effectiveness of these security metrics.
IV-A System Overview
Our proposed system follows the three-element model of Hendrikx et al. (section II-C) comprising a trustor, a trustee, and a trust engine. In the open-source supply chain, the trustor and trustee already exist—e.g., the maintainer team serves as the trustor, and a potential contributor is the trustee. Our work focuses on operationalizing the trust engine component, which is currently absent from the open-source ecosystem.
We describe our proposed system and case study in Figure 1. Interaction history defines the core of our reputation computation, and in our system, we focus specifically on security-related interactions defined by recommended security practices. Although some work has been done on trust establishment in OSS [51], the proposed frameworks are based on defining trust, our work operationalizes trust with reputation systems. In the next section, we define metrics to assess security interactions and history within a typical OSS ecosystem.
Table I: Proposed security signals. Bolded signal categories are derived from security recommendations (e.g, NIST, CNCF, etc). The proposed evaluation metrics in Table II offer ways to evaluate the signals below.
S/N. Proposed Signals & Analysis Metrics Description
SECURITY SIGNALS
1. Security vulnerability in artifact – Pull Requests/Commits  [51, 52, 53] Measures vulnerabilities introduced through pull requests or commits. Introduced either by the user or to a project owned by the user.
2. Security vulnerability in artifact – use of vulnerable dependencies  [51, 52] Assesses the use and introduction of vulnerable dependencies into the artifact.
3. Use of ecosystem code scanning and security analysis features[54] Evaluates usage of ecosystem’s security tools for code scanning.
4. Use of ecosystem integrity guarantees [27] Examines use of ecosystem integrity features (e.g., code signing).
5. Use of branch protection [55, 6] Checks if branch protection is enforced to prevent unauthorized changes.
6. Use of security policies and vulnerability reporting [55] Determines if the project has security policies and reporting mechanisms.
7. Use of automated workflows [55] Assesses use of [55] automation in workflows to enforce security.
REPUTATION WEIGHTAGE SIGNALS
W1. Package Usage[56] Weighs the utility of the potential contributor’s owned projects
W2. Community Tenure [51] Weighs the the length of the contributor’s belonging to the ecosystem
W3. Centrality Score[18] Weighs the degree of connections of the potential contributor to other actors in the ecosystem.
IV-B Interaction Data – Security Signals and Metrics Definitions
Our system computes reputation from an actor’s historical interactions within the ecosystem. We categorize these interactions by security signals and quantify each signal using measurable metrics.
To define appropriate security signals and metrics, we consider: (1) alignment with widely accepted security recommendations, (2) the actor’s demonstrated adherence to good security practices in previous contributions and to significant projects, and (3) a history of non-malicious contributions.
From these considerations, we derive our security metrics from two kinds of sources:
  1. 1.
    Security standards and recommendations: We consulted frameworks like the SLSA security framework [57], the CNCF software supply chain security guidelines [58], the NIST SSDF security framework [59], NIST SP 800-204D [60], Openssf S2C2F [61], and the CIS Software Supply Chain Security Guide[62]. Common recommendations across these sources were used to ensure the selection of well-established security practices.
  2. 2.
    Available security tools in the OSS ecosystem: We identified security tools available through GitHub’s user interface and API, which reflect the security capabilities easily available to contributors on the platform.
We grouped the resulting metrics into seven categories – Security Signals, based on contributors’ security tool usage and vulnerability management practices. A summary appears in  Table I. This resulted in seven security signals
IV-C Trust Engine – Reputation Computation
In this section, we outline the core functionality of the trust engine’s reputation score computation (see Figure 1). First, the interaction data formatter and security signal scorer extracts each contributor’s ecosystem interaction history and quantify each signal according to the metrics in Table II. These base measurements capture a user’s adherence to the defined security signals and their contribution patterns.
Next, ARMS evaluates each contributor’s reputation. To account for the risk posed by a contributor’s activities within the ecosystem, we refine the initial signal scores by incorporating:
Impact Score – (Weightage Signals Table I (W1-W3)):
  1. 1.
    Package Usage Weighting: Projects with no users (e.g., private repositories or those with zero forks/downloads) should not contribute to the user’s evaluation score, as they present minimal supply chain risk.
  2. 2.
    Community Tenure Weighting  [51]: This ensures that newer users do not unduly influence their reputation score, addressing concerns like those in the ”XZ utils” situation [63], where malicious activity was introduced by new members.
  3. 3.
    Centrality Score Weighting [18]: Centrality—an actor’s level of connectedness within the ecosystem based on contribution activities—can be crucial for scaling reputation. We suggest evaluating not only the number of connections (edges) but also the time taken to establish them.
Benchmark Score:
  1. 1.
    Combination of multiple security metrics/Composite Scoring: A holistic evaluation will aggregate a user’s performance across various security metrics, offering a comprehensive view of their adherence to security practices. This approach minimizes false positives and negatives by drawing from multiple metrics, as seen in the OpenSSF scorecard  [64]. Each metric’s weighting will be based on its effectiveness analysis section V-A.
  2. 2.
    Ecosystem Benchmarks: User performance will be benchmarked against the average or median reputation scores across the ecosystem to provide a relative measure of security adherence.
V Design of Experiments for ARMS
In this section, we outline potential studies to establish and evaluate the feasibility of an actor reputation system, exemplified by ARMS, and general use of actor metrics to establish trust in software supply chain security. These complementary studies—primarily observational or retrospective—aim to validate and refine the proposed security-reputation metrics and assess the real-world impacts of deploying ARMS. We categorize these studies into two groups: (1) evaluating the effectiveness of the proposed security metrics, and (2) examining user behaviors.
V-A Security Metrics Effectiveness Evaluation – Quasi-Experimental Analyses
To evaluate the utility of the proposed security signals, we propose the following quasi-experimental studies to determine whether the signals reliably predict poor security practices.
  1. 1.
    Inter-metric Relationship Study: Compare OSS projects that adopted a given security practice (e.g., branch protection) to matched control projects that did not, using a difference‐in‐differences design. We will analyze the results with regression models incorporating project and time fixed effects and interaction terms for the paired metrics to isolate their combined impact on security outcomes.
    • Data: Historical GitHub records for projects within the same domain, similar size, and activity levels.
    • Outcomes: Changes in vulnerability incidence (Signals 1–2) and shifts in contributor behaviors (Signals 3–7).
    • Ethics: Uses only publicly available data, avoiding additional data collection or privacy concerns.
  2. 2.
    Retrospective Incident Prediction Study: Identify projects that experienced known supply-chain incidents (e.g., npm/Event-Stream, ua-parser-js) and compare their pre-incident signal profiles against similar “clean” projects.
    • Data: Metric values for each project’s maintainers before the incident.
    • Analysis: Logistic regression (with interaction terms) to determine which signals best predict incident occurrence.
V-B User Behavioural Studies – Surveys & Interviews
To evaluate the practical usability of actor metrics in OSS Supply chain security, we propose the following studies:
  1. 1.
    Collaborator Vetting Study: We propose a Recruitment of active OSS maintainers to participate in a vignette study[65]. This would present anonymized contributor profiles with varying metric scores, and measure time‐to‐decision (vet vs. reject) and how variations on the proposed ARMS signals would influence these choices. This method would gather qualitative feedback on signal clarity and usefulness.
  2. 2.
    Chilling Effect: A potential issue with establishing an actor reputation system is the possibility of “chilling effect,”[66] where contributors may hesitate or reduce participation if they know their interactions are being tracked and scored [66, 67, 68]. We propose a user survey study to assess contributors’ willingness to participate under different tracking scenarios (e.g., “all projects” vs. “critical only”), and to quantify perceived privacy risks and impact of monitoring activities on contribution intent.
    Ultimately, we do expect some chilling effects. To mitigate them, we recommend limiting the deployment of an ARMS approach to high-impact, security-sensitive OSS projects (e.g., the Linux kernel) rather than applying it to hobby or low-risk repositories. To effectively do this, there is need to develop an effective project importance score. OpenSSF’s criticality score [69] may be used or serve as a starting point. We recommend surveying OSS contributors to gauge a suitable threshold of criticality.
V-C Worked Examples
We illustrate how our proposed system could have worked in the earlier described examples section III-B.
V-C1 XZ Utils Attack
The timeline of events leading to the XZ Utils backdoor reveals several characteristics that map to our proposed security signals [70]:
  1. 1.
    Recent account: The attacker created a GitHub account in January 2021 and joined the XZ Utils project in October 2021—well within their first year of activity.
  2. 2.
    Limited public history: Prior to October 2021, their contributions were confined to private repositories.
  3. 3.
    Targeted feature change requests: Their first public change request focused on adding features to a small set of projects rather than fixing issues.
Under our framework, these traits would yield a low reputation score:
  • Signal Metrics (Signals 1-7) penalize sparse or opaque contribution histories.
  • Community tenure (Signal W2) reduces scores for newly created accounts.
  • Centrality (Signal W3) remains low because the user’s contribution network is both recent and shallow.
V-C2 Dexcom
The Dexcom engineers’ case revealed repeated failures that lasted for extended periods. In one incident, an issue persisted from November 28 to December 3, 2019, halting Dexcom systems and severely impacting availability. Under our framework, this would be reflected in lower scores for Signals 1 and 2 (time to fix vulnerabilities and severity levels), as recurrent downtime directly reduces the reputation of engineers responsible for critical systems.
V-C3 ESLint Compromise
The 2018 ESLint compromise was traced to an attacker breaching a maintainer’s account—enabled by the absence of two-factor authentication. This type of attack cannot be modeled by reputation systems because the attacker took over a legitimate account without performing any genuine contributor actions or exhibiting the behavioral signals that these systems track. This is why we deem it out of scope in the system threat Model section III.
VI Discussions and Future Works
VI-A Threats to Validity
We begin by outlining some potential issues with our proposal:
  1. 1.
    False Positives and Negatives: There is a risk of mis-assigning reputation, resulting in inappropriate characterizations of users as more or less trustworthy. Our definitions and metrics do not account for all interactions that could detect incompetence or malicious activities, especially non-artifact interactions such as social engagements, organizations, security communications, feedback to code reviews, etc. To mitigate this, we recommend a weighted combination of metrics, with each metric’s effectiveness considered in section V-A. Using ecosystem-wide averages and standard deviations can also help raise the threshold for issuing advisories based on user scores.
  2. 2.
    Insider Threats: As a special case of false negatives, we emphasize that any reputational system is ineffective for insider threats where trust is deliberately built over time. However, such attacks are costly for an adversary.
  3. 3.
    Defining Security Interactions: As stated, our work envisions operationalizing actor trust security metrics; however, we acknowledge the inadequacies of our proposed security metrics. We encourage further research in defining trust, especially concerning the interactions between security metrics.
  4. 4.
    Privacy Concerns: Making scores public could unfairly harm honest actors. To address this, we propose that scores remain private, accessible on a need-to-know basis. Advisories based on these scores would be shared only with maintainers of projects to which the user wishes to contribute.
  5. 5.
    Gameability of Proposed Metrics: Our security metrics, like any trust-based system, can be exploited. Users may act, individually or in collusion, to enhance their reputation. We mitigate this risk with time-weighted scoring and recommend incorporating human oversight (e.g., reporters and moderators) for nuanced judgments. However, we acknowledge that trust and safety issues plague all online platforms [71], and a reputation system would be no exception.
  6. 6.
    Possibility of Chilling Effect: Discussed in section V-B.
VI-B Evaluating Actor Intent
Our work assumes that actors are well-intentioned and attributes security failures to genuine contributors’ lack of expertise or mistakes; it does not account for malicious intent. However, some supply-chain incidents arise from malicious actors using techniques such as account spoofing or credential theft. Incorporating intent into reputation systems substantially increases their complexity. Although our current security signals focus on expertise and behavior, they do not capture actor intent. Future work should extend these metrics to infer intent— for example, by analyzing anomalous contribution patterns, unusual social graph connections, or timing irregularities, ultimately creating a more comprehensive reputation model.
VI-C Supporting Ecosystem Heterogeneity
Reputation systems fundamentally face actor‐identification challenges [72, 73]. This issue is especially pronounced in open-source ecosystems, where contributors manage artifacts across multiple platforms—from source repositories to package registries. Although next-generation software-signing tools have improved artifact-to-actor verification [4, 74, 75], reliable cross-platform identification remains vulnerable to identity theft, impersonation, and other attacks. Consequently, ARMS requires stronger mechanisms to verify and vouch for identities across diverse environments.
To provide possible solutions, recent initiatives (e.g., CISA’s RFI for software identifier ecosystems [76]) propose establishing identities through multiple institutions—some designed to preserve privacy [77, 78]. Adapting these models across ecosystems introduces the dual challenges of federation—enabling actors in one ecosystem to be recognized in another—and roaming—allowing actors to transfer their identifier and reputation between providers. Federation and Roaming are challenges reminiscent to other federation protocols work (e.g., OIDC [79] and Distributed Identities [80]). Thus, for an actor reputation system like ARMS, institutions across ecosystems must collaborate and share reputation information whenever a software artifact from one ecosystem A is used in another ecosystem B.
VI-D Social & Process Signals: Beyond Artifact Contributions
Our current security (and weightage) signals (Table I) focus exclusively on an actor’s artifact contributions. As noted in our approach criticisms, we have not yet captured an actor’s interactions with other users—such as code-review feedback, issue discussions, and peer endorsements—which could provide valuable reputational insights.
We also recommend developing process-based metrics to evaluate informal OSS practices. Process-based methods [81] are most effective in structured development environments. However, in open-source ecosystems—where formal processes are often absent [82]—proxy process metrics can still gauge adherence to security best practices. For example, one could measure the frequency of explicit threat-model or design-review discussions in issues or change requests, track the presence and pass rate of automated security scans (e.g., SAST, SBOM generation) in continuous integration workflows, and monitor the regularity of dependency updates following published vulnerability disclosures. By combining artifact, social, and process signals, ARMS can deliver a more holistic reputation assessment for contributors in open-source ecosystems.
VI-E Ethical and Privacy Considerations
A system that records actor activities inevitably raises ethical and legal concerns. Actors must provide informed consent before their ecosystem activities are disclosed to projects they wish to join. At the same time, project owners need actionable insights into a contributor’s security posture. Balancing these interests requires privacy-preserving disclosure.
We propose that (i) contributors opt in to sharing reputational data, (ii) only aggregated or pseudonymized metrics (e.g., differential-privacy noise or percentile rankings) are revealed to maintainers, and (iii) contributors receive dashboards that show exactly what information will be shared. Compliance with regulations such as GDPR [83] should guide data-retention periods, user-deletion requests, and audit logging. Future prototypes should incorporate cryptographic approaches—such as zero-knowledge proofs or secure multiparty computation—to prove adherence to security metrics without exposing raw activity logs. Key directions include: Designing and evaluating privacy-preserving reputation protocols, conducting user studies to gauge contributor consent thresholds and maintainer information needs, and integrating regulatory compliance checks and automated audit trails into the prototype.
VI-F Why Focus Reputation on Cybersecurity?
We have situated the ARMS approach within the context of cybersecurity. The ARMS metrics can, of course, be extended in order to infer properties of engineers other than their cybersecurity expertise. We suggest that doing so for functional properties (i.e., input-output behaviors that can be validated through testing) may be unnecessary — standard engineering practices call for code contributions to be accompanied by adequate tests, such that functional properties can be assured through reference to the test results. Not so for non-functional properties, cybersecurity as one of many. It is for such properties that reputational measures may be more useful. For example, Cramer et al. recently reported that trust and safety defect repairs in social media platforms are rarely accompanied by automated tests, in part because the validation is apparently performed through use case analysis rather than through software behavioral analysis [71]. Similarly, behaviors for regulatory compliance such as GDPR are challenging to validate [11]. Since the current state-of-the-art software validation techniques cannot automatically conclude whether a system provides non-functional properties such as security, trust-and-safety, or regulatory compliance — at least not in a cost-effective manner — we think that ARMS approaches may be a suitable complementary measure of correctness.
VII Conclusion
In this paper, we explored the characteristics of the open-source software supply chain that necessitate actor-based security techniques. We advocate for the implementation of an actor reputation system to operationalize a framework for trust within the open-source ecosystem. We introduced seven security metrics for measuring reputation, alongside two evaluation methods: one for assessing actor performance and another for determining the effectiveness of these metrics in predicting reputation. Lastly, we identified future research directions, including empirical studies, system prototypes, and initiatives to maintain actor identities across ecosystems.
Appendix A Signal Metrics Definition Table
Table II: Proposed security signals, weighting factors, and evaluation metrics. Bolded security and weighting signals are derived from established frameworks (e.g., NIST, SLSA, CNCF) and prior work, respectively. The proposed metrics describe how to measure each highlighted security or weighting signal. Relevant case studies and related research are cited to justify inclusion of select metrics.
S/N. Proposed Signals & Analysis Metrics Description
SECURITY SIGNALS
1. Security vulnerability in artifact – Pull Requests/Commits Measures vulnerabilities introduced through pull requests or commits. Introduced either by the user or to a project owned by the user.
a. Time to fix/close security vulnerabilities of various severity levels [51] Time taken to address vulnerabilities detected in pull requests or commits.
b. Severity levels of security vulnerability reported [52] Evaluates the criticality of reported vulnerabilities (low, medium, high), for user’s projects.
c. Presence of vulnerability in PR/commits (not linked to external dependencies) [53] Measures severity of vulnerabilities that stem from user’s direct contributions.
2. Security vulnerability in artifact – use of vulnerable dependencies Assesses the use and introduction of vulnerable dependencies into the artifact.
a. Length of time before fix [51] Time taken to resolve vulnerable dependencies, after they have been detected (or publicized).
b. Severity levels of package’s reported vulnerability [52] Tracks severity of vulnerabilities in dependencies.
c. Number of Repos/Projects that have a vulnerable dependency Total number of projects affected by vulnerable dependencies as a percentage of users total projects.
d. Number of vulnerable dependencies/project Assesses how widespread vulnerable dependencies are within each project.
e. Number of projects with reported vulnerable dependencies Measures the overall exposure of projects to vulnerable dependencies.
3. Use of ecosystem code scanning and security analysis features Evaluates usage of ecosystem’s security tools for code scanning.
a. Status of dependabot alerts (for dependencies)[54] Monitors if security alerts for dependencies are being addressed.
b. Status of Secret sharing/Validity Checks Evaluates the management of secret-sharing mechanisms and validity checks.
c. Code scanning and Vulnerability alerts[54] Assesses if security scanning tools are actively used and vulnerability alerts managed.
d. Push Protection Status Checks if protection features are used to block pushes with exposed secrets (e.g., unscanned code).
4. Use of ecosystem integrity guarantees Examines use of ecosystem integrity features (e.g., code signing).
a. Number of Projects with an integrity guarantee [27] Measures how many projects are secured with integrity verification features such as code signing.
5. Use of branch protection Checks if branch protection is enforced to prevent unauthorized changes.
a. Number of protected branches per project Evaluates how many branches within each project are safeguarded against direct commits.
b. Number of Projects with Protected branches [55, 64] Tracks how many projects enforce branch protection.
6. Use of security policies and vulnerability reporting Determines if the project has security policies and reporting mechanisms.
a. Status of private vulnerability reporting, or security policy [55] Checks if private vulnerability reporting channels and security policies are established and functional.
7. Use of automated workflows Assesses use of [55] automation in workflows to enforce security.
a. Number of Projects with Automated workflows [55] Measures how many user’s projects use automated workflows for security enforcement, such as auto-deployments or auto-tests.
REPUTATION WEIGHTAGE SIGNALS
W1. Package Usage[56]
a. Number of downloads/stars/forks Weighs the potential contributor’s ownership of useful projects to the ecosystem.
W2. Community Tenure [51]
a. Length of time contributing to project Weighs how much time user has been a contributor to other projects.
b. Length of time owning account Weighs how long user has actually owned an account
c. Contributory Strength Weighs the strength of contributions to other projects by lines of code, issues, pull requests
W3. Centrality Score[18]
a. Connections to other Actors[18] Weighs a potential contributor’s network of interactions/contributory activities.
b. Time to form connections Length of time taken to form the connections in ’a’ above.